OpenStudy (russianboss1):

Help with math please fan and medal Which expression shows the solution of the system? {5x + 4y = 2 -2y = 8 + 4x A) |2 4| |5 2| |4 8 | |-2 4| --------------------- |5 4| |5 4| |-2 8||-2 8| B) |5 2| |2 4| |-2 4 | |4 8| --------------------- |5 4| |5 4| |-2 8||-2 8| C) A) |2 4| |5 2| |8 -2 | |-4 8| --------------------- |5 4| |5 4| |-4 -2||-4 -2| D) A) |5 2| |2 4| |-4 8 | |8 -2| --------------------- |5 4| |5 4| |-4 -2||-4 -2|

OpenStudy (russianboss1):

@ganeshie8

OpenStudy (russianboss1):

@AloneS

OpenStudy (texaschic101):

5x + 4y = 2 -4x -2y = 8...(2) -------------- 5x + 4y = 2 -8x -4y = 16 ------------- -3x = 18.....x = -6 5x + 4y = 2 5(-6) + 4y = 2 -30 + 4y = 2 4y = 32 y = 8

alones (alones):

Yo @RussianBoss1 may you post them one by one.. not just in one post.. cuz i get confused

OpenStudy (russianboss1):

you mean the answer choices?

OpenStudy (russianboss1):



OpenStudy (russianboss1):

@scoobydo

OpenStudy (scoobydo):

D it think